The state education commission wants to estimate the fraction of tenth grade students that have reading skills at or below the eighth grade level.

Step 2 of 2 :  

Suppose a sample of 2404 tenth graders is drawn. Of the students sampled, 1803 read above the eighth grade level. Using the data, construct the 98% confidence interval for the population proportion of tenth graders reading at or below the eighth grade level. Round your answers to three decimal places.
Answer #1

We have given,              
              
x=2404-1803=601          
n=2404          
Estimate for sample proportion=   0.25  
Level of significance is =1-0.98=0.02      
Z critical value(using Z table)=2.326      
              
Confidence interval formula is              


=(0.2295,0.2705)
              
Lower limit for confidence interval=0.230  
              
Upper limit for confidence interval=0.271